# Hard View Selection for Contrastive Learning

### v1

@@ 5.3 Q3: What are the Effects of Empowering the Adversary?
In this experiment, we briefly explore adversarially learning a network to output the transformation matrix for view generation. We optimize a Spatial Transformer Network (STN) (Jaderberg et al., 2015) jointly with the DINO objective and a ViT-tiny. Using the inverse of the DINO loss, we train the STN to output the geometric transformation matrices (each 6 parameters) for generating views given an input image that are used for standard CL training with DINO. We employ and test various transformation types (e.g., rotation, translation, etc.) that influence the capacity of the adversary. We discuss this experiment in greater detail in Appendix E and briefly summarize our findings here: applying an adversary in this way is very challenging due to the sensitivity of CL on the data augmentation and the dynamics of adversarial learning, often leading to either minor improvements or model collapses as we have observed when employing a too strong manual IoU policy in Q1.
@@
Figure 5: Comparing DINO with DINO+HVS on CIFAR100 for longer trainings.
Longer Training
Despite the inherent overhead associated with additional forward passes of HVS, our results clearly show that using HVS is also beneficial for long trainings (see 300-epoch improvements in Table 1). To delve deeper into the behavior of longer trainings under our computational constraints, we conduct several training runs with progressively higher epochs on the CIFAR100 dataset, using the same initial learning rate and a cosine schedule. The comparative analysis in Fig. 5 shows, except for some overfitting, no evidence of diminishing returns as the improvement remains consistent throughout training nor signs of instability (e.g., collapse) due to hard tasks.
@@ 6 Conclusion
In this study, we presented HVS, a new data augmentation and learning strategy for contrastive learning designed to challenge pretrained models with more demanding samples. This method, while straightforward in its design, has proven to be a powerful tool, pushing the boundaries of the traditional random view generation in contrastive learning. When combined with methods like DINO, SimSiam, and SimCLR, HVS consistently showcased improvements in linear evaluation and a diverse set of transfer tasks. As the landscape of self-supervised learning continues to evolve, methods like HVS underscore the potential for further advancements in the field. By continuously challenging models with harder samples and refining augmentation techniques, we anticipate even greater strides in model performance and generalization in future research endeavors. While HVS incurs a computational overhead due to its additional forward passes, our preliminary experiments indicate no signs of diminishing returns with extended training durations. This suggests that HVS is likely to continue to outperform baselines regardless of the training length. Looking ahead, we believe there is ample room for refining and expanding HVS, such as further investigating the number of sampled views based on the current learning state, enhancing efficiency through checkpointing activations, bypassing forwarding of similar pairs, or augmentation distributions that allow controlling task difficulty.
@@ G.3 Object Detection and Instance Segmentation
Appendix H Hard View Selection Objectives
H.1 SimCLR
@@
In this formulation, we iterate over all possible pairs of augmented images derived from the same original image across all ”N” augmentations, and compute the loss for each pair. The pair that gives the highest loss is then used in the optimization process. It is important to note that the temperature parameter $\tau$ and the similarity function remain as defined in the original loss function.

### v4

@@ 5.3 Q3: What are the Effects of Empowering the Adversary?
In this experiment, we briefly explore adversarially learning a network to output the transformation matrix for view generation. We optimize a Spatial Transformer Network (STN) (Jaderberg et al., 2015) jointly with the DINO objective and a ViT-tiny. Using the inverse of the DINO loss, we train the STN to output the geometric transformation matrices (each 6 parameters) for generating views . We employ and test various transformation types (e.g., rotation, translation, etc.) that influence the capacity of the adversary. We discuss this experiment in greater detail in Appendix E and briefly summarize our findings here: applying an adversary in this way is very challenging due to the sensitivity of CL on the data augmentation and the dynamics of adversarial learning, often leading to either minor improvements or model collapses as we have observed when employing a too strong manual IoU policy in Q1.
@@ 6 Conclusion
In this study, we presented HVS, a new data augmentation and learning strategy for Self-Supervised Learning designed to challenge pretrained models with more demanding samples. This method, while straightforward in its design, has proven to be a powerful tool, pushing the boundaries of the traditional random view generation in SSL. When combined with methods like DINO, SimSiam, iBOT, and SimCLR, HVS consistently showcased improvements in linear evaluation and a diverse set of transfer tasks. As the landscape of self-supervised learning continues to evolve, methods like HVS underscore the potential for further advancements in the field. By continuously challenging models with harder samples and refining augmentation techniques, we anticipate even greater strides in model performance and generalization in future research endeavors. While HVS introduces computational overhead due to its additional forward passes, our findings demonstrate that this overhead is significantly mitigated when using small resolution images in HVS without impeding its performance. This suggests that HVS is likely to continue to outperform baselines regardless of the training length.
@@ G.3 Object Detection and Instance Segmentation
Appendix H Computational Overhead of HVS
In Table 14, we compare the computational overhead factors across multiple baselines. The details of hardware and software used for this analysis are: NVIDIA RTX 3080, AMD R7 5800X, 32GB RAM, Ubuntu 22.04, PyTorch 2.0.1, CUDA 11.8. For DINO’s 2 global and 8 local views (default), applying HVS with nviews=2 sampled for each original view results in 4 global and 16 local views. Since considering all combinations would yield over 77k unique comparisons (4 over 2 times 16 over 8), to remain tractable, we limit the number of total comparisons to 128. Overall, reducing the image resolution can significantly reduce the time complexity of HVS. For instance, in the case of DINO ViT-S/16, the reduction leads to approximately 30% faster training times and as reported in the main paper, a 50% reduction allows retaining original performance.
Res. SimSiam (RN50) DINO (ViT)
100% x1.55 x2.15
75% x1.45 x1.68
50% x1.24 x1.56
25% x1.16 x1.5
Table 14: Time slowdown factors under different image resolutions for the selection in HVS.
While technically there can be a memory overhead with HVS, with the number of sampled views chosen in this paper, the backward pass of the methods that compute gradients only for the selected view pair still consumes more memory than the selection part of HVS (even for 8 sampled views in SimSiam). Note, that selection and the backward computation are never executed at the same time but sequentially.
We emphasize that the time overhead factors were measured without any optimization of HVS’ efficiency and, in our view, there are multiple ways to improve it. In our work, we opted for the easiest implementation possible to showcase that selecting harder views dependent on the model learning state can help boost performance in contrastive learning. Going forward with more compute-efficient HVS solutions, one could think of:
  • using embeddings of views from “earlier” layers in the networks or
  • using 4/8 bit low-precision for the view selection or
  • using one GPU just for creating embeddings and selecting the hardest views while the remaining GPUs are used for learning or
  • switching between HVS and the standard pipepline in alternating fashion or
  • bypassing forwarding of similar pairs and more
Appendix I Hard View Selection Objectives
I.1 SimCLR
@@
For each iteration, we evaluate all possible view pairs and contrast each view against every other example in the mini-batch. Intuitively, the pair that yields the highest loss is selected, which is the pair that at the same time minimizes the numerator and maximizes the denominator in the above equation. In other words, the hardest pair is the one, that has the lowest similarity with another augmented view of itself and the lowest dissimilarity with all other examples.
Appendix J Attention Maps
In this section, we visualize attention maps as a way to qualitatively study the potential effects and differences of the HVS learning method and their resulting features. The provided attention maps are from a DINO ViT-S/16 100 model, and we contrast the attention for each HVS and the baseline. All input images are from the ImageNet-1k validation split. The color code used depicts strong attention in yellow and weak attention in green and attentions from the HVS model are shown in the top row and attentions from the baseline in the bottom row, respectively. To summrize this study, we do not see apparent strong differences between HVS and the baseline. What we occasionally notice is that the HVS models seem to capture the shape of some subtle/indistinct objects better (e.g. the creek/river in the valley in Fig. 12, the lizard in Fig. 13, or the speaker in Fig. 14 or focus slightly more on the context and background (Fig. 16, 17, and 18).
Figure 11:
Figure 12:
Figure 13:
Figure 14:
Figure 15:
Figure 16:
Figure 17:
Figure 18:
Figure 19: